OpenStudy (anonymous):

? The mean and standard deviation of a population being sampled are 64 and 6, respectively. If the sample size is 50, what is the standard error of the mean? how do I find the standard error?

OpenStudy (kropot72):

The standard error of the sample mean is: \[\frac{\sigma}{\sqrt{n}}=\frac{6}{\sqrt{50}}\]

OpenStudy (anonymous):

so the answer is .85?

OpenStudy (kropot72):

Correct.

OpenStudy (anonymous):

Thank you! would you be able to help me with my next one? @kropot72

OpenStudy (kropot72):

Please post it :)

OpenStudy (anonymous):

In a random sample of 85 people attending a workshop, 63 of the respondents said the workshop was useful. What is the sample proportion of the people who said the workshop was useful? would the answer be .13?

OpenStudy (kropot72):

The sample proportion is: \[\frac{63}{85}=you\ can\ calculate\]

OpenStudy (anonymous):

Thank you! I got it:)

OpenStudy (kropot72):

Good work :)